OpenStudy (anonymous):

GIven: AB is a tangent of point X. AB=70 and XD=12. What is the measure of DB

OpenStudy (anonymous):

Diagram

OpenStudy (raden):

u mean the length of DB ?

OpenStudy (anonymous):

Yes, what is the length of DB

OpenStudy (raden):

if yes, use the pythagorean theorem : DB^2 = AD^2 + AB^2 with AD = 2DX

OpenStudy (anonymous):

So the answer would be 74, correct?

OpenStudy (raden):

yes, i got like that too

OpenStudy (anonymous):

Okay can you help me with one more?

OpenStudy (anonymous):

Given: DC is a tangent of E. AE = 20 and DC = 9 what is the length of AC?

OpenStudy (raden):

do same idea like 1st, AC^2 = AD^2 + CD^2 (with AD = 2AE)

OpenStudy (anonymous):

So 41? Thanks!!

OpenStudy (raden):

yep, that's right